# The Directionality of Gravitational and Thermal Diffusive Transport in Geologic Fluid Storage

### v2

@@ Abstract
Diffusive transport has implications for the long-term status of underground storage of hydrogen (H2) fuel and carbon dioxide (CO2), technologies which are being pursued to mitigate climate change and advance the energy transition. Once injected underground, CO2 and H2 will exist in multiphase fluid-water-rock systems. The partially-soluble injected fluids can flow through the porous rock in a connected plume, become disconnected and trapped as ganglia surrounded by groundwater within the storage rock pore space, and also dissolve and migrate through the aqueous phase once dissolved. Recent analyses have focused on the concentration gradients induced by differing capillary pressure between fluid ganglia which can drive diffusive transport (“Ostwald ripening”). However, studies have neglected or excessively simplified important factors; namely: the non-ideality of gases under geologic conditions, the opposing equilibrium state of dissolved CO2 and H2 driven by the partial molar density of dissolved solutes, and entropic and thermodiffusive effects resulting from geothermal gradients. We conduct an analysis from thermodynamic first principles and use this to provide numerical estimates for CO2 and H2 at conditions relevant to underground storage reservoirs. We show that while diffusive transport in isothermal systems is upwards for both gases, as indicated by previous analysis, entropic contributions to the free energy are so significant as to cause a reversal in the direction of diffusive transport in systems with geothermal gradients. For CO2, even geothermal gradients less than 10oC/km (far less than typical gradients of 25oC/km) are sufficient to induce downwards diffusion at depths relevant to storage. Diffusive transport of H2 is less affected, but still reverses direction under typical gradients, e.g. 30oC/km, at a depth of 1000 m. This reversal occurs independent of the solute’s thermophobicity or thermophilicity in aqueous solutions. The entropic contribution also modifies the magnitude of flux where geothermal gradients are present, with the largest diffusive fluxes estimated for CO2 with a 30oC/km gradient, despite the higher diffusion coefficient of H2. We find a maximum flux on the order of 10-13 mol/cm2/s for CO2 in the 30oC/km scenario; significantly lower than literature estimates for maximum convective fluxes in moderate to high permeability formations. Contrary to previous studies, we find that in diffusion and convection will likely work in concert – both driving CO2 downwards, and both driving H2 upwards – for conditions representative of their respective storage reservoirs.
@@ II.2 Gas Ganglia
With fugacity calculated from the EoSs noted above, we use eqn. 15 to determine the partitioning relationship between concentration and fugacity of the pure gas phase over our considered range of depths (3). Fugacity coefficient is defined as the ratio of fugacity to system pressure, $\phi=\frac{f}{P}$. Partition coefficients (ratios quantifying how a species will be distributed in two phases at equilibrium) calculated from eq. 15 are presented as gas phase fugacity to aqueous solubility, with units of $\frac{MPa}{\text{mol gas / mol aq. phase}}$, simplified to MPa.
@@ III Impact of Temperature Gradients
In order to calculate the impact of thermal gradients, We obtain values for molar entropy $S_{m}$ for CO2 and H2 under our considered pressure and temperature conditions via the following steps. The molar entropy for an ideal gas ($S^{ig}$) at a given temperature ($S^{ig}_{T}$)is found using the Shomate equation with parameters tabulated by [51], calculated from data originally from Chase [62], and modified for pressure through $S^{ig}=S^{ig}_{T}-Rln(P/P_{R})$ ($P_{R}$ is the reduced pressure $P/P_{c}$ where $P_{c}$ is the pressure at the critical point). Deviation from the ideal gas value is found using a departure function [63, 64], calculated using the compressibility factor and constants calculated in the Peng-Robinson EOS [32]. Under our considered conditions, CO2 molar entropy ranges from approximately 140-220 J/mol-K, and H2 from 80-110 J/mol-K; these calculated values are consistent with tabulated data [65] to approx. 2%. Higher pressure decreases the molar entropy while higher temperature increases molar entropy; the relationship of entropy to depth is thus nontrivial and dependent on geothermal gradient (Figure 5).
@@ IV Kinetics of Diffusion
As shown in Figure 8, diffusive flux in the isothermal case is always positive (diffusion transports molecules upwards), and estimated values for CO2 and H2 are similar (in the region relevant to storage of both gases), on the order of $10^{-15}-10^{-14}$ mol/cm2/s. As observed above, flux quickly becomes negative (diffusive transport is downwards) for CO2 under even the low geothermal gradient of 10oC/km; H2 shows a more subtle dependence, only achieving negative fluxes below 1000 m for the 30oC/km case. In the geothermal gradient case (30oC), downwards CO2 flux is several times larger than the flux of H2; its magnitude is an order larger than the isothermal CO2 case, on the order of $10^{-13}$ mol/cm2/s.
@@ IV.1 Comparison with Mass Transport via Convective Dissolution
To provide an order-of-magnitude comparison, we refer to Neufeld et al. [13] who estimated a convective CO2 flux equalling 20 kg m-2 yr-1 for the Sleipner site in the North sea, which is equivalent to 1.44$\times$10-9 mol CO2/cm2/s. Compared to the upper limit of the diffusive CO2 flux estimated above, the convective flux is $\approx 4$ orders of magnitude larger (depending on temperature assumption). The estimate of Neufeld et al. [13] represents an upper bound for storage reservoirs, due to the high permeability (5,000 mD per Michael et al. [67]) of the storage reservoir at Sleipner. Emami-Meybodi et al. [14] show that the maximum of time-resolved dissolution flux for CO2 (i.e. prior to convective shutdown) decreases by $\approx$ 2 orders of magnitude when Rayleigh number (a dimensionless number describing convection, proportional to permeability) is decreased by that same factor. This indicates that even for the maximal assumptions considered above, diffusive transport of CO2 will be a minor factor in moderate to high-permeability formations. Li et al. [24] provide a more thorough analysis of convective transport rates for a range of scenarios and similarly conclude that convective transport is likely to initiate and complete well before diffusive transport becomes significant, for all scenarios considered, except in low permeability ($\leq$ 1 mD) formations. For both diffusion and convection, flux is time-dependent, and this simple order-of-magnitude comparison of the maximal cases does not apply for all time periods. In particular, the low-permeability case where both convection and diffusion act on similar timescales and with similar fluxes requires further study to determine the combined effect on transport. However, it is important to note that in situations with positive geothermal gradients (i.e. all likely storage reservoirs), we predict that diffusion will drive CO2 downwards, thus increasing storage security even in the low permeability case – in the geothermal case, diffusion and convection of CO2 are not acting in competing directions.
@@ V Conclusions
We provide an upper estimate for flux rates, using assumptions favorable to faster diffusion. The maximum diffusive flux estimated for CO2 occurs under our largest investigated geothermal gradient of 30oC and is on the order of 10-13 mol/cm2/s; four orders of magnitude smaller than the downwards convective flux estimated for CO2 under high-permeability conditions by Neufeld et al. [13]. We note that under low permeability conditions and for later times, the fluxes of convection and diffusion may become comparable. Regardless, our analysis indicates that for all typical CO2 storage scenarios, regardless of permeability, both diffusion and convection act in concert to drive CO2 downwards.

### v3

@@ Abstract
Diffusive transport has implications for the long-term status of underground storage of hydrogen (H2) fuel and carbon dioxide (CO2), technologies which are being pursued to mitigate climate change and advance the energy transition. Once injected underground, CO2 and H2 will exist in multiphase fluid-water-rock systems. The partially-soluble injected fluids can flow through the porous rock in a connected plume, become disconnected and trapped as ganglia surrounded by groundwater within the storage rock pore space, and also dissolve and migrate through the aqueous phase once dissolved. Recent analyses have focused on the concentration gradients induced by differing capillary pressure between fluid ganglia which can drive diffusive transport (“Ostwald ripening”). However, studies have neglected or excessively simplified important factors; namely: the non-ideality of gases under geologic conditions, the opposing equilibrium state of dissolved CO2 and H2 driven by the partial molar density of dissolved solutes, and entropic and thermodiffusive effects resulting from geothermal gradients. We conduct an analysis from thermodynamic first principles and use this to provide numerical estimates for CO2 and H2 at conditions relevant to underground storage reservoirs. We show that while diffusive transport in isothermal systems is upwards for both gases, as indicated by previous analysis, entropic contributions to the free energy are so significant as to cause a reversal in the direction of diffusive transport in systems with geothermal gradients. For CO2, even geothermal gradients less than 10oC/km (far less than typical gradients of 25oC/km) are sufficient to induce downwards diffusion at depths relevant to storage. Diffusive transport of H2 is less affected, but still reverses direction under typical gradients, e.g. 30oC/km, at a depth of 1000 m. This reversal occurs independent of the solute’s thermophobicity or thermophilicity in aqueous solutions. The entropic contribution also modifies the magnitude of flux where geothermal gradients are present, with the largest diffusive fluxes estimated for CO2 with a 30oC/km gradient, despite the higher diffusion coefficient of H2. We find a maximum flux on the order of 10-13 mol/(cm2s) for CO2 in the 30oC/km scenario; significantly lower than literature estimates for maximum convective fluxes in moderate to high permeability formations. Contrary to previous studies, we find that in diffusion and convection will likely work in concert – both driving CO2 downwards, and both driving H2 upwards – for conditions representative of their respective storage reservoirs.
@@ II.2 Gas Ganglia
With fugacity calculated from the EoSs noted above, we use eqn. 15 to determine the partitioning relationship between concentration and fugacity of the pure gas phase over our considered range of depths (3). Fugacity coefficient is defined as the ratio of fugacity to system pressure, $\phi=\frac{f}{P}$. Partition coefficients (ratios quantifying how a species will be distributed in two phases at equilibrium) calculated from eq. 15 are presented as gas phase fugacity to aqueous solubility, with units of MPa/(mol gas/mol aq. phase), simplified to MPa.
@@ III Impact of Temperature Gradients
In order to calculate the impact of thermal gradients, We obtain values for molar entropy $S_{m}$ for CO2 and H2 under our considered pressure and temperature conditions via the following steps. The molar entropy for an ideal gas ($S^{ig}$) at a given temperature ($S^{ig}_{T}$)is found using the Shomate equation with parameters tabulated by [51], calculated from data originally from Chase [62], and modified for pressure through $S^{ig}=S^{ig}_{T}-Rln(P/P_{R})$ ($P_{R}$ is the reduced pressure $P/P_{c}$ where $P_{c}$ is the pressure at the critical point). Deviation from the ideal gas value is found using a departure function [63, 64], calculated using the compressibility factor and constants calculated in the Peng-Robinson EOS [32]. Under our considered conditions, CO2 molar entropy ranges from approximately 140-220 J/(mol-K), and H2 from 80-110 J/(mol-K); these calculated values are consistent with tabulated data [65] to approx. 2%. Higher pressure decreases the molar entropy while higher temperature increases molar entropy; the relationship of entropy to depth is thus nontrivial and dependent on geothermal gradient (Figure 5).
@@ IV Kinetics of Diffusion
As shown in Figure 8, diffusive flux in the isothermal case is always positive (diffusion transports molecules upwards), and estimated values for CO2 and H2 are similar (in the region relevant to storage of both gases), on the order of $10^{-15}-10^{-14}$ mol/(cm2s). As observed above, flux quickly becomes negative (diffusive transport is downwards) for CO2 under even the low geothermal gradient of 10oC/km; H2 shows a more subtle dependence, only achieving negative fluxes below 1000 m for the 30oC/km case. In the geothermal gradient case (30oC), downwards CO2 flux is several times larger than the flux of H2; its magnitude is an order larger than the isothermal CO2 case, on the order of $10^{-13}$ mol/(cm2s).
This simplified analysis provides estimates of maximum flux, only reflective of the hypothetical ganglia-initialized state. Further work is needed to establish more comprehensive time-resolved kinetics, including the impacts of density-driven convection and thermal diffusive processes (e.g. Haugen and Firoozabadi [67]).
@@ IV.1 Comparison with Mass Transport via Convective Dissolution
To provide an order-of-magnitude comparison, we refer to Neufeld et al. [13] who estimated a convective CO2 flux equalling 20 kg m-2 yr-1 for the Sleipner site in the North sea, which is equivalent to 1.44$\times$10-9 mol CO2/(cm2s). Compared to the upper limit of the diffusive CO2 flux estimated above, the convective flux is $\approx 4$ orders of magnitude larger (depending on temperature assumption). The estimate of Neufeld et al. [13] represents an upper bound for storage reservoirs, due to the high permeability (5,000 mD per Michael et al. [68]) of the storage reservoir at Sleipner. Emami-Meybodi et al. [14] show that the maximum of time-resolved dissolution flux for CO2 (i.e. prior to convective shutdown) decreases by $\approx$ 2 orders of magnitude when Rayleigh number (a dimensionless number describing convection, proportional to permeability) is decreased by that same factor. This indicates that even for the maximal assumptions considered above, diffusive transport of CO2 will be a minor factor in moderate to high-permeability formations. Li et al. [24] provide a more thorough analysis of convective transport rates for a range of scenarios and similarly conclude that convective transport is likely to initiate and complete well before diffusive transport becomes significant, for all scenarios considered, except in low permeability ($\leq$ 1 mD) formations. For both diffusion and convection, flux is time-dependent, and this simple order-of-magnitude comparison of the maximal cases does not apply for all time periods. In particular, the low-permeability case where both convection and diffusion act on similar timescales and with similar fluxes requires further study to determine the combined effect on transport. However, it is important to note that in situations with positive geothermal gradients (i.e. all likely storage reservoirs), we predict that diffusion will drive CO2 downwards, thus increasing storage security even in the low permeability case – in the geothermal case, diffusion and convection of CO2 are not acting in competing directions.
@@ V Conclusions
We provide an upper estimate for flux rates, using assumptions favorable to faster diffusion. The maximum diffusive flux estimated for CO2 occurs under our largest investigated geothermal gradient of 30oC and is on the order of 10-13 mol/(cm2s); four orders of magnitude smaller than the downwards convective flux estimated for CO2 under high-permeability conditions by Neufeld et al. [13]. We note that under low permeability conditions and for later times, the fluxes of convection and diffusion may become comparable. Regardless, our analysis indicates that for all typical CO2 storage scenarios, regardless of permeability, both diffusion and convection act in concert to drive CO2 downwards.